If one-half of one integer is subtracted from three-fifths of the next consecutive integer, the difference is 3. What are the two integers?
Answer by jmg(22)

It talks about an integer and the next consecutive integer
So you would let the integer be n, and the next consective would be 1 more than that number so it would be (n+1)
Using the question:
If one-half of one integer is SUBTRACTED FROM three-fifths of the next, the difference is 3.
Just change the words to numbers and symbols.
1/2(n) is SUBTRACTED FROM 3/5(n+1) = 3 , so you would actually write it
3/5(n + 1) - 1/2(n) = 3 Distribute
3/5(n) + 3/5 - 1/2(n) = 3 Combine your like terms the 3/5 and 1/2 n
1/10(n) + 3/5 = 3 Move the 3/5 to the other side
1/10(n) = 12/5 Multiply by the reciprocal of 1/10, which would be 10/1 in order to eliminate the fraction in front of the n.
This gives you n = 24, which is the integer, so the next consecutive integer would be 24 + 1 = 25.
So your numbers are 24 and 25
Hope this helps